# Federated Foundation Model for Cardiac CT Imaging

### v1

@@ Abstract
Federated learning (FL) is a renowned technique for utilizing decentralized data while preserving privacy. However, real-world applications often involve inherent challenges such as partially labeled datasets, where not all clients possess expert annotations of all labels of interest, leaving large portions of unlabeled data unused. In this study, we conduct the largest federated cardiac CT imaging analysis to date, focusing on partially labeled datasets ($n=8,124$) of Transcatheter Aortic Valve Implantation (TAVI) patients over eight hospital clients. Transformer architectures, which are the major building blocks of current foundation models, have shown superior performance when trained on larger cohorts than traditional CNNs. However, when trained on small task-specific labeled sample sizes, it is currently not feasible to exploit their underlying attention mechanism for improved performance. Therefore, we developed a two-stage semi-supervised learning strategy that distills knowledge from several task-specific CNNs (landmark detection and segmentation of calcification) into a single transformer model by utilizing large amounts of unlabeled data typically residing unused in hospitals to mitigate these issues. This method not only improves the predictive accuracy and generalizability of transformer-based architectures but also facilitates the simultaneous learning of all partial labels within a single transformer model across the federation. Additionally, we show that our transformer-based model extracts more meaningful features for further downstream tasks than the UNet-based one by only training the last layer to also solve segmentation of coronary arteries. We make the code and weights of the final model openly available, which can serve as a foundation model for further research in cardiac CT imaging.
1 Introduction
@@
The manual annotation of medical images is a laborious task that requires expert knowledge [41, 38]. Often, physicians can only label a limited amount of data for deep learning model training. They typically focus on labeling data relevant to their specific research needs, leaving a significant portion of data unlabeled and thus unused for training. As a result, small, highly specialized subsets of large, mostly unlabeled datasets are common in local clinics. This presents two opportunities for improvement. First, the training data can be enlarged by leveraging all labeled subsets across clinics, while accounting for the different structures annotated in each. Second, by leveraging labeled and unlabeled datasets in a pooled training synergy effects can be realized, if over all participating hospitals every label of interest is present in at least one location. Additionally, the diversity of training data from various institutions can expand the overall training distribution (Fig. 1).
@@
Privacy laws hinder the widespread collection of such heterogeneous large scale datasets stored at a single location [^1] Health Insurance Portability and Accountability Act (HIPAA): https://www.cdc.gov/phlp/publications/topic/hipaa.html, General Data Protection Regulation (GDPR): https://gdpr.eu/. Federated Learning (FL) is one paradigm that circumvents privacy concerns by reverting the paradigm of central data storage [40, 18, 17, 42]. In FL, the model is distributed to all data holding institutions, where training is performed locally before the model is sent back to a central server. On this server the trained model weights from all participating clients are averaged before another round of training is initialized (see Fig. 1). Unfortunately, the quality and consistency of labels across different clients can vary, impacting the model’s performance. Without inspection from the data scientist label quality and consistency must be ensured in FL, which often poses a big challenge that impedes the predictive performance of federated trained models on real world data [18].
@@
In situations where each institution has a different subset of the total training labels, the clients are termed partially labeled. Training on such clients requires complex algorithms for handling the loss computation on clients, where labels are not present. Partially labeled data can further result in a skewed distribution of labels across clients. Some labels might be overrepresented in the overall dataset, while others are underrepresented. This can lead to biased models that perform well on some labels of data but poorly on others. Training a single model to effectively address all tasks across these clients is challenging due to the uneven distribution of annotations.
@@
The largest FL study on 3D medical images to date ($n=6,314$) was performed by [35], who trained an automatic tumor boundary detector for the rare disease of glioblastoma in a federated manner. They reported improvement over a publicly trained model especially on rare cases that are not represented in rather small public datasets. Other works include the prediction of future oxygen requirement of COVID 19 patients and the histological response to breast cancer [9, 49]. The largest federated learning study in cardiovascular imaging is conducted by Linardos et al. [24]. They use subsets of the publicly available magnetic resonance imaging (MRI) datasets from the Multi-Centre, Multi-Vendor, and Multi-Disease Cardiac Image Segmentation Challenge (M&Ms) and Automated Cardiac Diagnosis Challenge (ACDC) with 180 patients in total [6, 5]. In all the aforementioned studies, it is assumed that all clients possess all labels available in the federation. However, the scenario of partially labeled clients has not been considered.
All approaches report an increase in generalizability for the federated trained model compared to the individual trained one. However, in all reported studies on real world data the labels were consistently distributed across all participating clients in the federation. To the best of our knowledge there exists no comparable study with federated learning on real world data with a similar amount of data on partially labeled clients. Additionally, unlabeled data is usually discarded and not used to further increase model performance. Compared to the segmentation of tumors on MR brain scans [35] our considered use case has a much larger variability in the used data.
When solving imaging tasks a convolutional UNet is most often the method of choice [35, 20, 15, 55]. Due to the inductive bias of the convolutional operations they tend to generalize better with smaller amounts of data than transformer based architectures [29]. However, recent literature on foundation models has shown promising performance of transformers when only trained on large enough dataset sizes [33, 51, 25]. Transformers have the advantage of a larger receptive field due to their inbuilt attention mechanism. Not exhibiting an inductive bias also becomes an advantage if only enough data samples are used. To improve the performance of both model architectures knowledge distillation (KD) can be used. In KD, usually a student network is trained to mimic output of a teacher network [14].
Diseases of the cardiovascular system amount for up to a third of deaths in developed countries [^2] https://www.who.int/news-room/fact-sheets/detail/cardiovascular-diseases-(cvds). A common valve pathology is described by aortic valve stenosis, which is a condition where the aortic valve becomes narrowed, leading to reduced blood flow from the heart to the rest of the body. A Transcatheter Aortic Valve Implantation (TAVI) is primarily performed catheter-based to replace the present valve with an artificial one. Due to its less invasive nature it has become the gold standard for treating severe aortic stenosis in patients who are considered high risk or inoperable for surgical aortic valve replacement [22, 11]. However, patients receiving a TAVI are more prone to be dependent on a pacemaker post implantation due to the prosthesis applying pressure to the stimulation conduction system of the heart [43]. Known influencing parameters are the aortic valve geometry, the per-cusp calcification, and the distance of the annulus plane to the membranous septum [28, 32]. The three hinge points determine the location of the aortic annulus plane, which is the location of the smallest diameter of the aortic root and thus determines the size of the prosthesis, while the coronary ostia determine the possible length. A measurement not yet taken in clinical practice is the location of the smallest part of the membranous septum and its distance from the annulus plane [16]. Multiple works exist that perform localization of aortic root and hinge points as well as coronary ostia [21, 3, 2]. However, all methods were trained on single site data and no method is able to quantify all aspects from the CT, which might be attributed to missing labels for some subtasks at the specific locations. Thus, the presented approaches might not generalize well outside of their training datasets. No method exists so far that combines quantification of aortic landmarks with the detection of the membranous septum and the quantification of calcification in the area of the aortic root, which are both important predictors for selecting the right prosthesis and determine the outcome after TAVI.
@@
Within the German Center for Cardiovascular Diseases (DZHK) we have set up a federated learning infrastructure connecting eight cardiology and radiology departments of university clinics in Germany. Each institution provides the pre-TAVI computed tomography (CT) scans as well as their individual present labels that help in choosing the right prosthesis (hinge points, coronary ostia, membranous septum, and per-cusp calcification). All CT scans used in the federated training are routinely acquired in clinical practice. One caveat of dealing with real world clinical data is the heterogeneity of available labels, which is especially prominent in our use case. While the annotated hinge points and coronary ostia are labeled across all participating locations, the membranous septum as well as the calcification are only labeled at a few not completely overlapping institutions. This work shows the following contributions.
@@
  • Study size and label scarcity: We present the largest up to date study in cardiac computed tomography imaging from real world patient data spanning eight hospitals in Germany ($n=8,124$). In our study, labels are scarce, meaning not all locations are in the possession of all label categories and further only a small fraction of data samples are labeled at the respective locations.
@@
  • Federated point detection and segmentation: We train a model for each custom task (hinge points and coronary ostia points, points of membranous septum, and segmentation of calcification) in a federated manner [15]. Due to their inductive bias convolutional neural networks can generalize better with small amounts of training samples. For guidance we feed the segmented heart as a condition such that the models can learn the anatomical relations between heart and the corresponding structures. We show the superiority of the federated approach for each subtask.
@@
  • Transfer into one model: We are the first to employ federated knowledge distillation to fuse the knowledge of the per-task models into a different architecture than the teacher, a transformer network when small amounts of manual annotations are available. The two-stage approach increases the amount of training data mitigating the performance difference between transformer and convolutional UNet by semi-supervised learning.
@@
  • Privacy-Preserving Label Quality Visualization: Due to its privacy by design structure FL does not enable the inspection of label quality at the participating sites. To verify consistency we compare the relative location of landmarks across locations, which does not disclose patient information but allows for qualitatice privacy-preserving outlier detection in label quality.
@@
  • Open source code and model weights: The code will be made publicly available. Further, we release the model weights of the final transformer model, which can be used as a foundation model in cardiac CT imaging for future studies.
Figure 1: Overview of federated consortium and federated knowledged distillation (KD) training pipeline. A Each label subset is not present at all locations. One model (UNet) is trained for each subset in a federated manner across the clients in possession of that label. B Subsequently, the federated trained models are used to make predictions on the unlabeled data samples. C The transformer based- model (SWIN-UNETR [12]) is trained from the predictions of the teacher network with three heads but the same backbone. Last, only the heads are fintuned on the human annotated data samples.
2 Results
@@
(Semi-supervised) Federated Knowledge Distillation from Partially labeled Clients
The large subset of unlabeled medical data usually remains unused when training deep learning models. On the other hand the most advanced deep learning architectures based on the transformer architecture can only develop their full potential when trained on large scale datasets. We opt for combining both worlds by first training a conventional UNet on the labeled subsets of all the data. By employing the federated approach for all models we were able to include heterogeneous data from multiple locations and thus widen the training distribution for each submodel. Last, we predicted on the large datasets that were unlabeled. Subsequently, we trained a transformer architecture (SWIN-UNETR) [12] to mimic the output from the UNets on the large datasets. We used the federated trained UNets from all three subtasks. Thus, we condense the knowledge for all three tasks from three models into one with semi-supervised federated KD. Due to the mutual exclusiveness of the labels, the model gained one head for each subtask while using the same backbone. With this design, the network has the freedom to learn features that are task-agnostic in their backbone, e.g. features for segmenting the aortic root and for localizing the three aortic valve hinge points are likely to be rather similar. After KD, each head is fine-tuned on the expert-annotated datasets across the consortium, while leaving the weights of the backbone feature extractor fixed.
@@
For evaluation purposes we performed a large series of experiments comparing different architectures and local vs. federated training. For each task (point detection of hinge points, coronary ostia, membranous septum, and segmentation of calcification) three different methods are compared. First, we train a convolutional UNet and transformer-based (SWIN-UNETR) on each local dataset. Second, both models are trained in a federated fashion across the locations having labels. Third, we performed semi-supervised federated knowledge distillation on the unlabeled data of each hospital with our federated trained UNet as the teacher and a SWIN-UNETR as student, before finetuning on the labeled subsets. For federated training we always left at least one location out for training for having an independent testset in form of a completely separated client.
Figure 2: Data distribution across clients. In total $8,124$ CT scans are available across all eight clients. For each label the distribution is differently skewed. While the most uniform distribution is present for the hinge point training, for membranous septum and calcification the distribution is skewed. Still, the federated model that is trained over these skewed distributions exhibits better performance than the one trained on a single client. HPs & COs: Hinge Points and Coronary Ostia Points, MS: Membranous Septum.
Figure 3: Qualitative results of the predicted labels of FedKD SWIN-UNETR, which is the final distilled model. The predictions were inspected by two experienced cardiologists verifying that the points are placed within the anatomical variance present. RCC: right coronary cusp, LCC: left coronary cusp, NCC: non-coronary cusp, RCO: right coronary ostium, LCO: left coronary ostium, MS1: upper, and MS2: lower point of membranous septum, Myo: myocardium, LA: left atrium, LV: left ventricle, RA: right atrium, RV: right ventricle, PA: pulmonary artery.
As can be seen in Fig. 4 models trained only on the local data shards underperform on datasets from other locations. Transformer based architectures generalize worse than convolutional UNet based ones, which we attribute to the inherent inductive bias of these architectures. The mean distance of the predicted hinge points of the local UNet approach is at $3.09\pm 1.71\mathrm{mm}$ for the same location and at $3.80\pm 2.02\mathrm{mm}$ for held out test locations, while the SWIN-UNETR predicts points at a mean distance of $2.60\pm 1.65\mathrm{mm}$ and $3.73\pm 1.97\mathrm{mm}$ respectively. Federated training improves generalization performance for both methods. However, the UNet ($2.59\pm 1.76\mathrm{mm}$, $3.43\pm 1.79\mathrm{mm}$) performs better than the transformer ($3.06\pm 1.70\mathrm{mm}$, $3.89\pm 1.91\mathrm{mm}$). While the performance of the SWIN-UNETR can be enhanced by performing semi-supervised federated knowledge distillation from the federated trained UNet on the previously unlabeled data samples at all locations the performance of the KD UNet is similar to the federated one. The predicted points lie at a mean distance of $2.80\pm 1.71\mathrm{mm}$ for the training locations and $3.36\pm 1.83\mathrm{mm}$ for the held out test locations for the transformer and at $3.18\pm 1.92\mathrm{mm}$ and $3.83\pm 2.12\mathrm{mm}$ for the UNet.
The performance for detecting the membranous septum is similar to localizing the hinge points. The UNet generalizes better with fewer data samples, but the SWIN-UNETR can be improved with semi-supervised federated knowledge distillation to even surpass the UNet on the unseen test clients. The local UNet predicts a mean distance of $3.01\pm 1.84\,\mathrm{mm}$ on the same client and $4.30\pm 1.82\,\mathrm{mm}$ on others, the local transformer predicts points at $3.96\pm 2.19\,\mathrm{mm}$ and $4.06\pm 2.16\,\mathrm{mm}$ distance respectively. When training both models in a federated manner the SWIN-UNETR generalizes better when knowledge distillation is employed, what is seen with the lower standard deviation. The UNet’s mean distance lies at $3.40\pm 1.56\,\mathrm{mm}$, while the transformer’s is at $3.29\pm 1.45\,\mathrm{mm}$. The performance on the training clients is very similar (UNet: $2.99\pm 1.81\,\mathrm{mm}$, KDT: $2.95\pm 1.72\,\mathrm{mm}$).
Segmenting the calcification in the aortic root leads to different results than the previous tasks. The transformer performs better than the UNet especially when trained in a federated manner with a DICE score of $68.27\pm 20.15$ on the testsets of the training locations and $69.15\pm 23.23$ on the held out test location. The federated UNet only achieves a DICE score of $49.45\pm 20.94$ and $39.10\pm 21.20$ respectively. The model trained with KD is almost on par with the federated trained transformer with DICE scores of $64.55\pm 20.79$ and $67.01\pm 23.14$. We attribute the slightly worse performance to the concurrent point detection which seems to favor partly other image features than calcification segmentation.
In conclusion, semi-supervised federated knowledge distillation enhances the predictive performance of a transformer based architecture (SWIN-UNETR) [12] to be better or on par with the Unet based counter part. Further, the tasks of locating the hinge points, coronary ostia, and membranous septum as well as segmentiing the calcification of the aortic root can be solved with one model despite the distributed label classes across different classes.

### v2

# Real World Federated Learning with a Knowledge Distilled Transformer for Cardiac CT Imaging

@@ Abstract
Federated learning is a renowned technique for utilizing decentralized data while preserving privacy. However, real-world applications often face challenges like partially labeled datasets, where only a few locations have certain expert annotations, leaving large portions of unlabeled data unused. Leveraging these could enhance transformer architectures’ ability in regimes with small and diversely annotated sets. We conduct the largest federated cardiac CT analysis to date ($n=8,104$) in a real-world setting across eight hospitals. Our two-step semi-supervised strategy distills knowledge from task-specific CNNs into a transformer. First, CNNs predict on unlabeled data per label type and then the transformer learns from these predictions with label-specific heads. This improves predictive accuracy and enables simultaneous learning of all partial labels across the federation, and outperforms UNet-based models in generalizability on downstream tasks. Code and model weights are made openly available for leveraging future cardiac CT analysis.
Introduction
@@
The manual annotation of medical images is a laborious task that requires expert knowledge [1, 2]. Often, physicians can only label a limited amount of data for deep learning model training. They typically focus on labeling data relevant to their specific research needs, leaving a significant portion of data unlabeled and thus unused for training. As a result, small, highly specialized subsets of large, mostly unlabeled datasets are common in local clinics. This presents two opportunities for improvement. First, the training data can be enlarged by leveraging all labeled subsets across clinics, while accounting for the different structures annotated in each. Second, by leveraging labeled and unlabeled datasets in a pooled training synergy effects can be realized, if over all participating hospitals every label of interest is present in at least one location. Additionally, the diversity of training data from various locations can expand the overall training distribution (Figure 1).
@@
Privacy laws hinder the widespread collection of such heterogeneous large scale datasets stored at a single location [3]. Federated Learning (FL) is one paradigm that circumvents privacy concerns by reverting the paradigm of central data storage [4, 5, 6, 7]. In FL, the model is distributed to all data holding locations, where training is performed locally before the model is sent back to a central server. On this server the trained model weights from all participating locations are averaged before another round of training is initialized (see Figure 1a). Unfortunately, the quality and consistency of labels across different locations can vary, impacting the model’s performance. Without inspection from the data scientist label quality and consistency must be ensured in FL, which often poses a big challenge that impedes the predictive performance of federated trained models on real world data [5].
@@
In situations where each hospital has a different subset of the total training labels, the locations are termed partially labeled. Training on such locations requires complex algorithms for handling the loss computation, where labels are not present. Partially labeled data can further result in a skewed distribution of labels across locations. Some labels might be overrepresented in the overall dataset, while others are underrepresented. This can lead to biased models that perform well on some labels of data but poorly on others. Training a single model to effectively address all tasks across these locations is challenging due to the uneven distribution of annotations.
@@
The largest FL study on 3D medical images to date ($n=6,314$ patients) was performed by Pati et al. [8], who trained an automatic tumor boundary detector for the rare disease of glioblastoma in a federated manner. They reported improvement over a publicly trained model especially on rare cases that are not represented in rather small public datasets. Other works include the prediction of future oxygen requirement of COVID 19 patients, the histological response to breast cancer, and the diagnosis of hypertrophic cardiomyopathy from ECG and Echocardiograms [9, 10, 11]. The largest federated learning study in 3D cardiovascular imaging is conducted by Linardos et al. [12]. They use subsets of the publicly available magnetic resonance imaging (MRI) datasets from the Multi-Centre, Multi-Vendor, and Multi-Disease Cardiac Image Segmentation Challenge (M&Ms) and Automated Cardiac Diagnosis Challenge (ACDC) with 180 patients in total [13, 14]. In all the aforementioned studies, it is assumed that all locations possess all labels available in the federation. All approaches report an increase in generalizability for the federated trained model compared to the individual trained one. To the best of our knowledge there exists no comparable study with federated learning on real world data on partially labeled datasets. Additionally, unlabeled data is usually discarded and not used to further increase model performance.
In this work, we present a solution to train federated deep learning networks when imaging labels are scarce and their distributions are highly imbalaced over many locations. This presents a scenario where recent transformer architectures have severe limitations due to their dependence on large labeled cohorts [15]. The key contribution of this work is to use techniques from knowledge distillation to substantially increase the performance of these architectures for the purpose of leveraging their strengths towards solving related downstream tasks on the same type of images. Transformers, with their inherent attention mechanism, benefit from a larger receptive field and the absence of inductive biases becomes advantageous in data-rich scenarios [48, 17].
Figure 1: Overview of federated consortium and federated knowledge distillation (KD) training pipeline. a) Federated learning procedure and b) our consortium across eight university hospitals in Germany. c) Each label subset is not present at all locations (Stage 1A). One model (UNet) is trained for each subset in a federated manner across the locations in possession of that label. d) Subsequently, the federated trained models are used to make predictions on the unlabeled data samples (Stage 1B). e) The transformer based- model is trained from the predictions of the teacher network with three heads but the same backbone (Stage 2AB). Last, only the heads are fintuned on the human annotated data samples. Naming is consistent with Figure 7.
When solving imaging tasks a convolutional UNet is most often the method of choice [18, 19, 8, 20]. Due to the inductive bias of the convolutional operations they tend to generalize better with smaller amounts of data than transformer based architectures [21]. To bridge the gap between architectures that excel in low-data regimes (e.g., convolutional networks) and those that require large datasets (e.g., transformers), knowledge distillation (KD) can be used. Three primary types of distillation can be distinguished: response-, feature-, and relation-based [22, 23]. Response-based KD focuses on aligning the final output predictions (logits) of the teacher and student models [24, 25]. Its architecture-agnostic nature makes it particularly suitable when teacher and student models differ in structure, as it trains the student to mimic the teacher’s output distributions for better generalization. Feature-based KD, on the other hand, aligns the intermediate feature representations of both models, requiring compatible architectures to effectively match features at specific layers [26, 27]. Finally, relation-based KD emphasizes teaching the student to capture relationships between data samples as modeled by the teacher [28, 29]. Originally proposed for model compression, where a smaller student network learns to mimic the outputs of a larger teacher network [24], KD extends well beyond this use case. All above KD approaches are at their core fundamentally a method for transferring knowledge from a teacher model to a student model, regardless of their sizes or architectures [30]. Additionally, recent works on foundation models have effectively utilized KD to transfer knowledge from large pre-trained teacher models to student models, highlighting its usability in training large scale models [15].
@@
Within the German Center for Cardiovascular Diseases (DZHK) we have set up a federated learning infrastructure connecting eight cardiology and radiology departments of university clinics in Germany. Each location provides CT scan of patients with aortic stenosis and the corresponding label types. One caveat of dealing with real world clinical data is the heterogeneity of available labels, which is especially prominent in our use case. While the annotated hinge points and coronary ostia are labeled across all participating locations, the membranous septum as well as the calcification are only labeled at a few not completely overlapping locations. Furthermore, a large quantity of CT data is completely unlabeled. Consequently, our approach includes two major factors that enlarge the data distribution used for model training: 1) the unlabeled data samples in the hospitals and 2) the federated learning approach. More precisely, this work shows the following contributions.
@@
  • Study size and label scarcity: We present the largest up to date study in cardiac computed tomography imaging from real world patient data spanning eight hospitals in Germany ($n=8,104$ images). In our study, labels are scarce, meaning not all locations are in the possession of all label categories and further only a small fraction of data samples are labeled at the respective locations.
@@
  • Federated point detection and segmentation: We train a convolutionalmodel for each custom task i.e. label type (hinge points and coronary ostia points, points of membranous septum, and segmentation of calcification) in a federated manner, which we refer to as stage 1 of our proposed learning method [18]. Due to their inductive bias convolutional neural networks can generalize better with small amounts of training samples. We show the superiority of the federated approach for each subtask.
@@
  • Semi-supervised two stage learning strategy: We are the first to employ federated knowledge distillation (stage 2) to fuse the knowledge of the per-task models (from stage 1) into a different architecture than the teacher when small amounts of manual annotations are available. With the CNNs predictions are generated on the unlabeled datasets functioning as pseudo labels for training the transformer-based architecture. The two-stage approach increases the amount of training data mitigating the performance difference between transformer and convolutional UNet by semi-supervised learning.
@@
  • Privacy-Preserving Label Quality Visualization: Due to its privacy by design structure FL does not enable the inspection of label quality at the participating sites. To verify consistency we compare the relative location of landmarks across locations, which does not disclose patient information but allows for qualitative privacy-preserving outlier detection.
@@
  • Open source code and model weights: The code will be made publicly available. Further, we release the model weights of the final transformer model, which can be used as a base model in cardiac CT imaging for future studies.
Our use case for the developed method focuses on improving the analysis of cardiac CT imaging for Transcatheter Aortic Valve Implantation (TAVI) patients. Diseases of the cardiovascular system amount for up to a third of deaths in developed countries [32]. A common valve pathology is described by aortic valve stenosis, which is a condition where the aortic valve becomes narrowed, leading to reduced blood flow from the heart to the rest of the body. TAVI is a catheter-based procedure to replace the narrowed valve with an artificial one, necessitating precise imaging and analysis for optimal outcomes. Due to its less invasive nature it has become the gold standard for treating severe aortic stenosis in patients who are considered high risk or inoperable for surgical aortic valve replacement [33, 34]. However, patients receiving TAVI are more prone to be dependent on a pacemaker post implantation due to the prosthesis applying pressure to the stimulation conduction system of the heart [35]. Known influencing parameters are the aortic valve geometry, the per-cusp calcification, and the distance of the annulus plane to the membranous septum [36, 37]. The three hinge points define the location of the aortic annulus plane, which is the location of the smallest diameter of the aortic root and, thus, determines the size of the prosthesis, while the coronary ostia determine the possible length. A measurement not yet taken in clinical practice in an automatic way is the location of the smallest part of the membranous septum and its distance from the annulus plane [38]. Multiple works exist that perform localization of aortic root and hinge points as well as coronary ostia [39, 40, 41]. All methods were trained on single-site data, lacking the ability to quantify all CT aspects due to missing labels for certain subtasks. Consequently, these approaches may not generalize well beyond their training datasets. No existing method combines aortic landmark detection with membranous septum detection and aortic root calcification quantification, which are key predictors for prosthesis selection and TAVI outcomes.
Results
The results are presented as follows. First, we describe how federated training enhances performance. Second, we present the effects of our proposed two-stage learning procedure. Next, we assess the consistency and reliability among labelers in a privacy-preserving manner by evaluating label quality using known anatomical relationships. This leads us to examine the impact of inter-observer variability on model predictions, a critical issue in federated learning. Finally, we evaluate our model’s generalization performance on a public dataset for a different task.
@@ Results
(Semi-supervised) Federated Knowledge Distillation from Partially Labeled Datasets
@@
For evaluation purposes we perform a large series of experiments comparing different architectures and local vs. federated training. For each task (point detection of hinge points, coronary ostia, membranous septum, and segmentation of calcification) three different methods are compared with different models. We train a convolutional UNet as well as two transformer architecture (ViT for segmentation and SWIN-UNETR) [42, 43]. While ViT is based on conventional self-attention, SWIN-UNETR employs a shifted window attention approach that trades of global and local context. First, we train a UNet and the two transformer-based models on each local dataset. Second, all models are trained in a federated fashion across the locations having these labels. Third, we perform semi-supervised federated knowledge distillation on the unlabeled data of each hospital with our federated trained UNet as the teacher and a ViT as well as a SWIN-UNETR as student, before finetuning on the labeled subsets. For federated training we always leave at least one location out for training for having an independent testset in form of a completely separated dataset.
The results are presented quantitatively in Table 1 (mean and standard deviation) and their distribution in Figure 2 (boxplot with median and quartiles) and qualitatively in Figure 3. Due to the obvious underperformance of ViT we do not present the results in Figure 2 to remain a direct comparison of the UNet and our best performing transformer-based architecture (SWIN-UNETR). The results for all tasks and models per location can be found in the supplementary information.